OpenStudy (anonymous):

Find the three arithmetic means between 5 and 17.

OpenStudy (anonymous):

6 7 and 8?

OpenStudy (anonymous):

not at all sure what the question is asking are you?

OpenStudy (anonymous):

nope lol

OpenStudy (anonymous):

then you can't do it

OpenStudy (er.mohd.amir):

Arithmetic mean of a and b is (a+b)/2

OpenStudy (anonymous):

8 11 14

OpenStudy (anonymous):

apparently the answer is 8, 11, 14 i have no idea why

OpenStudy (anonymous):

thats what i thought thanks

OpenStudy (anonymous):

those are evenly spaced for sure, but the question, as i read it, makes no sense at all

OpenStudy (er.mohd.amir):

let x1 ,x2, x3 be te three AM between 5 and 17 then total terms is n=5 first term is 5 and common difference is d and last term is 17 then l=a+(n-1)d 17=5+(5-1)d d=3 x1=a+d=5+3=8 x2=a+2d=11 x3=a+3d=14